Given the linear equation y=4x-2
Give the line, in slope-intercept form, that is perpendicular to this line and passes through the point (12,7)

Answer by mananth(16946)

1 y = 4 x -2
Divide by 1
y = 4 x + -2
Compare this equation with y=mx+b, m= slope & b= y intercept
slope m = 4

The slope of a line perpendicular to the above line will be the negative reciprocal -0.25
m1*m2=-1
The slope of the required line will be - 1/4

m= - 1/4 ,point ( 3 , 7 )
Find b by plugging the values of m & the point in
y=mx+b
7 = -3/4 + b
b= 31/4
m= - 1/4
The required equation is ==>y =- 1/ 4x+31/4
m.ananth@hotmail.ca